×
×
RSS Feed

IP-ONLY LANDVETTER / HÄRRYDA

Scheduled on 2017-10-17 09:00:00 Status Resolved Fault / Issue Estimated finish 2017-10-17 14:00:00

Påverkat område: Härryda/ Landvetter.
Påverkade postnummer: 438 93

Orsak till driftstörning: Strömavbrott.

Aktuell lägesrapport:

2017-10-17 09:13 - Larm om att vi har tappat kontakten med access switch in kommer, Felsökning påbörjad.
2017-10-17 09:20 - Lokalt strömavbrott som beräknas vara klart 14:00

Related servers / services

IP-ONLY ÅRE

Scheduled on 2017-10-17 00:00:00 Status Finished Estimated finish 2017-10-17 00:00:00

Påverkat område: Åre, Sweden
Påverkade postnummer: 830 13, 830 15, 837 98

Orsak till driftstörning: Okänt

Aktuell lägesrapport:
2017-10-17 00:22 - Larm indikerar problem.
2017-10-17 01:00 - Felsökning pågår.
2017-10-17 01:50 - Tekniker bokad.
2017-10-17 06:30 - Tekniker på väg till nod.
2017-10-17 07:30 - Tekniker på väg till nod.
2017-10-17 08:30 - Tekniker på plats och börjar felsöka.
2017-10-17 09:30 - Felsökning pågår.

Related servers / services

IP-ONLY FIBER - AVGRÄVD KABEL - PÅVERKAR ENBART NORRA NORRLAND

Scheduled on 2017-09-28 15:00:00 Status Finished Estimated finish 2017-09-29 06:00:00

Incident start: 2017-09-28 15:04
Incident description: Network disturbance due to Third party cable cut
Current estimated time of repair/resolution(ETR): No ETR

Affected geographical area: Stockholm - Norrland


SWE: internet-trafiken går en annan väg men om möjligt kan det vara lite skakigt och även gå trögt. Tekniker håller på att svetsa ihop den trasiga fiber nu.

Impact: Outage
Magnitude: low
Cause of disturbance: Third party cable cut

Current progress report:

2017-09-28 15:04 - Alarms raised, troubleshooting initiated.
2017-09-28 16:15 - Confirmed with supplier that a third party have a on going disturbance. Technicians en route.
2017-09-28 17:00 - Technicians en route
2017-09-28 18:00 - Technicians en route
2017-09-28 19:00 - Technicians on site and measuring the fibre.
2017-09-28 20:00 - Technicians have located the cable cut, it is exaclty in the edge of the road.
2017-09-28 20:30 - They have ordered a truck so they can re route the traffic on the road. They estimate that they can begin working on the fibre 2017-09-29 00:00
2017-09-28 22:30 - Awaiting arrival of TMA truck.
2017-09-29 06:00 - Sub provider reports that all circuits should be restored and technicians are hands off.

Related servers / services